OpenStudy (scarlettthecamel):

The volume of a piece of drainpipe is 25 litres. If the diameter of the pipe is 5cm, how long is it? Give your answer in metres to 2 decimal places. (Remember, 1 litre = 1000cl, 1cl = 1cm^3)

jigglypuff314 (jigglypuff314):

hi, do you still need help with this?

OpenStudy (scarlettthecamel):

yes @jigglypuff314

jigglypuff314 (jigglypuff314):

well if the diameter is 5 then the radius would be 2.5 right? :)

OpenStudy (scarlettthecamel):

yes i have had other people say that but i don't get how the radius has to do with it

OpenStudy (scarlettthecamel):

is it becoz of pi r^2

jigglypuff314 (jigglypuff314):

yes :) volume of pipe = pi r^2 L

OpenStudy (scarlettthecamel):

okay so how do i use this to get to the length

jigglypuff314 (jigglypuff314):

from that equation L = V / (pi r^2) where V = 25 * 1000 and r = 2.5

OpenStudy (scarlettthecamel):

okay

jigglypuff314 (jigglypuff314):

so L = (25000)/ (pi (2.5)^2)

OpenStudy (scarlettthecamel):

thanks so what does '/' mean?

jigglypuff314 (jigglypuff314):

oh sorry xD it means division, divide

OpenStudy (scarlettthecamel):

oright thanks haha!

OpenStudy (scarlettthecamel):

so would it be 1273.2

OpenStudy (scarlettthecamel):

@jigglypuff314

jigglypuff314 (jigglypuff314):

yes, that's what I got :) sorry for the late reply